EXERCISE 10–4 Direct Labor and Variable Manufacturing Overhead Variances LO10–2, LO10–3 Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows: Standard Hours Standard Rate per Hour Standard Cost 18 minutes $17.00 $5.10 During August, 5,750 hours of direct labor time were needed to make 20,000 units of the Jogging Mate. The direct labor cost totaled $102,350 for the month. Required: What is the standard labor-hours allowed (SH) to makes 20,000 Jogging Mates? What is the standard labor cost allowed (SH × SR) to make 20,000 Jogging Mates? What is the labor spending variance? What is the labor rate variance and the labor efficiency variance? The budgeted variable manufacturing overhead rate is $4 per direct labor-hour. During August, the company incurred $21,850 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month.
Answer #1

1) Standard labor hour allowed = 20000*18/60 = 6000 Hours

2) Standard labor cost allowed = 6000*17 = $102000

3) Labor spending variance = 102000-102350 = 350 U

4) Labor rate variance = (17*5750-102350) = 4600 U

Labor efficiency variance = (6000-5750)*17 = 4250 F

5) Variable overhead rate variance = (4*5750-21850) = 1150 F

Variable overhead efficiency variance = (6000-5750)*4 = 1000 F